OpenStudy (anonymous):

Solve the system y = -x + 7 and y = 0.5(x - 3)2

OpenStudy (anonymous):

y = 0.5(x - 3)2 -> y = x - 3

OpenStudy (anonymous):

Can you confirm that 2y = 4

OpenStudy (anonymous):

x=26/3, y=-5/3

OpenStudy (anonymous):

\[\{x=5,y=2\} \]

OpenStudy (triciaal):

y = -x + 7 and y = 0.5(x - 3)2 -x + 7 = (1/2)x^2 -1/2(6x) +1/2(9) multiply everything by the lcd 2 the rearrange x^2 - 4 x -5 = 0 (x - 5)(x + 1) = 0 x = 5, x = -1 y = -x + 7 when x = 5: -5 + 7 = 2 when x = -1 y = -(-1) + 7 = 8 (5, 2) and (-1, 8) check in y = 0.5(x - 3)2 left =2 right 0.5(5-3)^2 = 0.5(2)^2= 2 left = right therefore (5,2) correct left 8 right 0.5(-1 - 3)^2 = 0.5(-4)^2 = 0.5(16) = 8 left = right therefore (-1, 8) correct